What is the expected after-tax cash flow from selling a piece of equipment if Litchfield Design purchases the equipment today for 130,000 dollars, the tax rate is 30 percent, the equipment is sold in 3 years for 15,000 dollars, and MACRS depreciation is used where the depreciation rates in years 1, 2, 3, 4, and 5 are 35 percent, 28 percent, 23 percent, 9 percent, and 5 percent, respectively?
Answer #1

Book value as on date of sale=Cost-Accumulated Depreciation

=130,000*(1-0.35-0.28-0.23)

=18200

Hence loss on sale=18200-15000=3200

Hence after-tax cash flow=Sale proceeds+(Tax rate*Loss on sale)

=15000+(3200*0.3)

=$15960